A constant torque of 35 Nm is applied to a wheel whose moment of inertia is 1.5 kg m2. Find the Angular Acceleration. Also, find the angular speed at the end of 25 seconds


1
Expert's answer
2021-11-14T17:10:42-0500

By the definition of the torque we have:


"\tau = I\alpha,""\alpha=\dfrac{\tau}{I}=\dfrac{35\ N\times m}{1.5\ kg\times m^2}=23.3\ \dfrac{rad}{s^2}."

We can find the angular speed at the end of 25 seconds as follows:


"\omega=\omega_0+\alpha t,""\omega=0+23.3\ \dfrac{rad}{s^2}\times25\ s=582.5\ \dfrac{rad}{s}."